Hi I need help on understanding this problem:
give exact and approximate solutions.
(x-5)^2=18

Answer by stanbon(75887)

give exact and approximate solutions.
(x-5)^2=18
---
Exact Solutions:
(x-5)^2 - 18 = 0
Factor:
(x-5 - sqrt(18))(x-5 + sqrt(18)) = 0
x = 5+3sqrt(2) or x = 5 - 3sqrt(2)
----------------------------------------
Approximate Solutions:
x = 5 + 4.2426.. or x = 5 - 4.2426..
---
x = 9.2426.. or x = 0.7574..
Cheers,
Stan H.